# Symmetric Entropy-Constrained Video Coding for Machines

### v1

@@ IV-F3 Ablation on the SPDF Module
To explore how to effectively utilize pixel-oriented features $f_{\textit{pixel}}$ for MVS-oriented reconstruction, channel-wise concatenation ($M_{8}$) and direct addition ($M_{9}$) are employed to replace the gated fusion, as shown in Table VI. Experimental results demonstrate that gated fusion outperforms the other two fusion methods. Notably, the rate-task performance drops significantly by $19.5\%$ when $f_{\textit{pixel}}$ is not used ($M_{10}$). Furthermore, visualization results indicate that the gated fusion mechanism can selectively combine $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$ for final reconstruction. As shown in Fig. 13, $f_{\textit{semantic}}$ plays the major role in the reconstruction on channel 9, while contributing only features related to boats, forests, lake surface, and sky on channels 21, 25, and 36.
@@ V Conclusion
In this paper, we propose a unified Symmetric Entropy-Constrained Video Coding framework for Machines (SEC-VCM). By aligning the video coding process with the video understanding process of the visual backbone from an entropy perspective, SEC-VCM eliminates machine-vision-irrelevant information. The proposed BiEC mechanism softly maintains consistency between the visual backbone’s encoding process and the video decoding process, while the semantic-pixel dual-path fusion module introduces pixel-level priors to suppress artifacts and enhance reconstruction quality. Extensive experiments on diverse downstream tasks demonstrate the superiority and generalization ability of our method in achieving MVS-friendly reconstruction. We hope our proposed method offers insights into bridging the gap between neural video coding and machine vision, serving as a step forward for advancing the field of video coding for machines (VCM).

### v2

@@ IV-F3 Ablation on the SPDF Module
To explore how to effectively utilize pixel-oriented features $f_{\textit{pixel}}$ for MVS-oriented reconstruction, channel-wise concatenation ($M_{9}$) and direct addition ($M_{10}$) are employed to replace the gated fusion, as shown in Table VI. Experimental results demonstrate that gated fusion outperforms the other two fusion methods. Notably, the rate-task performance drops significantly by $19.5\%$ when $f_{\textit{pixel}}$ is not used ($M_{11}$). Furthermore, visualization results indicate that the gated fusion mechanism can selectively combine $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$ for final reconstruction. As shown in Fig. 13, $f_{\textit{semantic}}$ plays the major role in the reconstruction on channel 9, while contributing only features related to boats, forests, lake surface, and sky on channels 21, 25, and 36.
@@ V Conclusion
In this paper, we propose a unified Symmetric Entropy-Constrained Video Coding framework for Machines (SEC-VCM). By aligning the video coding process with the video understanding process of the visual backbone from an entropy perspective, SEC-VCM eliminates machine-vision-irrelevant information. The proposed BiEC mechanism softly maintains consistency between the visual backbone’s encoding process and the video decoding process, while the semantic-pixel dual-path fusion module introduces pixel-level priors to suppress artifacts and improve downstream task performance. Extensive experiments on diverse downstream tasks demonstrate the superiority and generalization ability of our method in achieving MVS-friendly reconstruction.
Nevertheless, our approach still has limitations. Developing a more lightweight framework and incorporating multimodal large language models into the downstream task evaluation will be directions for future work. We believe our proposed method offers insights into bridging the gap between neural video coding and machine vision, serving as a step forward for advancing the field of video coding for machines.